OpenStudy (anonymous):

The statement "cot =12/5 , sec = -13/5, and the terminal point determined by is in quadrant 4": A. cannot be true because 12^2+5x^2 not equal to 1 . B. cannot be true because cot must be less than 1. C. cannot be true because if cot =-12/5 , then sec = +_13/12. D. cannot be true because sec is greater than zero in quadrant 4.

OpenStudy (anonymous):

sorry dont know this one.

OpenStudy (anonymous):

its okay, thanks for trying to help :)

OpenStudy (luigi0210):

Here this will help

OpenStudy (anonymous):

what do i need to do with the graph?

OpenStudy (luigi0210):

It shows where each trig function is positive right? Look at your statement carefully

OpenStudy (anonymous):

it cant be true because if cot =-12/5 , then sec = +_13/1 2 this statement is correct.

OpenStudy (anonymous):

i mean incorrect.

OpenStudy (anonymous):

im going with C , thanks @Luigi0210

OpenStudy (luigi0210):

@Jhannybean finish it?

OpenStudy (jhannybean):

No thanks. Go ahead.

OpenStudy (luigi0210):

No, I'm a bit occupied at the moment.. please? >.<

OpenStudy (anonymous):

at quadrant IV cos supposed to be +ve and Cot also should be -ve

OpenStudy (anonymous):

since any negative is < 1, cot must not be >1 at IV quadrant. we can't determine cot at IV quadrant because it is greater than 1.

OpenStudy (anonymous):

hey i got it wrong i gueesed lol. thanks though!

OpenStudy (anonymous):

u can use tan and cot identity formula that equals to 1

OpenStudy (anonymous):

sin^2 θ + cos^2 θ = 1 tan^2 θ + 1 = sec^2 θ cot^2 θ + 1 = csc^2 θ these are Pythagorean Identities. Fundamental trig identitiy

OpenStudy (anonymous):

@StudentLearner it okay.

OpenStudy (anonymous):

d